# Aleph-Alpha-GermanWeb: Improving German-language LLM pre-training with model-based data curation and synthetic data generation

### v2

@@ Abstract
Scaling data quantity is essential for large language models (LLMs), yet recent findings show that data quality can significantly boost performance and training efficiency. We introduce a German-language dataset curation pipeline that combines heuristic and model-based filtering techniques with synthetic data generation. We use our pipeline to create Aleph-Alpha-GermanWeb, a large-scale German pre-training dataset which draws from: (1) Common Crawl web data, (2) FineWeb2, and (3) synthetically-generated data conditioned on actual, organic web data. We evaluate our dataset by pre-training both a 1B Llama-style model and an 8B tokenizer-free hierarchical autoregressive transformer (HAT). A comparison on German-language benchmarks, including MMMLU, shows significant performance gains of Aleph-Alpha-GermanWeb over FineWeb2 alone. This advantage holds at the 8B scale even when FineWeb2 is enriched by human-curated high-quality data sources such as Wikipedia. Our findings support the growing body of evidence that model-based data curation and synthetic data generation can significantly enhance LLM pre-training datasets.
@@ 1 Introduction
In recent years, ever-larger and more capable large language models (LLMs) have been released. This trend has led to the identification of power-law correlations between the loss value and the number of LLM parameters or pre-training dataset size [20, 47]. As a result, researchers have argued [20] that simply increasing the number of LLM parameters without also increasing the pre-training dataset size yields diminishing returns. Additionally, considering the cost of LLM inference [42] in production use-cases – which can far exceed the model’s initial training cost – practitioners may weigh the cost-benefit trade-off of replacing a larger model with a smaller one to minimise inference costs [22]. To compensate for the reduced model size, a sufficiently-large training data budget would be required [10, 17]. However, this approach poses significant challenges for domains and languages where data is scarce.
Many recent examples in both pre-training [29, 46] and post-training [36, 53] of LLMs indicate that improvements to data quality can make training much more efficient. This is because higher-quality data can reduce the data quantity needed to reach or exceed the same model performance. In fact, conventional data scaling laws, which suggest that a small percentage increase in performance requires increasing the amount of training data by an order of magnitude, appear to be broken when data quality is improved [45]. As a result, improving data quality allows us to train LLMs much more economically. Indeed, a recent 1.7 billion parameter language model pre-trained on 11 trillion tokens achieved new state-of-the-art results for its size by focussing heavily on data quality [3].
@@
  1. a)
    removing syntactic (near) duplication, e.g., identical documents (or with slightly different wording), which are typically straightforward to identify and remove [24];
  2. b)
    removing semantic duplication, e.g., different summaries of the same book, which we expect to maintain some of for the purposes of syntactic variance, but should probably only keep in proportion to the level of importance or complexity of the underlying concept [1];
  3. c)
    removing bad data, e.g., code that doesn’t compile [12] or language which is ungrammatical [38], less our models learn to repeat these mistakes; and
  4. d)
    augmenting or upsampling good data, e.g., adding high-quality synthetic data generated by LLMs [27] or intentionally duplicating high-quality subsets of the data [21].
Many curation methods originated as human-led and heuristic approaches, e.g., preselecting sources with a reputation for high quality data or manual annotation and filtering of datasets [39, 38]. More recently, however, model-based approaches, e.g., training annotation models to classify data into different quality buckets [25] or using LLM-based text quality filtering [46], have been shown to outperform heuristic methods. Additionally augmenting datasets with synthetically-generated data conditioned on high-quality organic data sources not only improves quantity but has been shown to further improve quality [46].
To date, model-based and synthetic approaches exist mostly for English-language text [25, 46]. One of the most commonly-used and largest heuristically-filtered non-English web datasets is FineWeb2 [21], which adapts prior heuristic methods shown to be successful for English web data [39, 38]. Recently, model-based methods have been applied to a subset of languages from FineWeb2 to score them for quality, finding as little as 15% of the original dataset can be used to achieve comparable performance on 1 billion parameter LLMs [31].
Nevertheless, there remains an overall quantity issue; FineWeb2, which includes data from over 1,000 languages, is <20% the size of the English-only FineWeb dataset [38] (8TB vs. 47TB in total disk size). A single language like German constitutes <1.4% of FineWeb2, and model-based approaches only reduce this quantity further [31]. Meanwhile, synthetic approaches to improve the quantity, diversity, and quality of non-English language pre-training data have focussed on machine translation [50], which comes with a host of socio-technical issues [32], and may lack the distinct syntactic or semantic advantages offered by synthetic approaches for LLM pre-training [28]. Here we aim to fill these gaps by constructing a dataset for state-of-the-art German pre-training.
@@
In this paper we present a data curation pipeline specifically for German-language LLM pre-training. We use it to derive our own dataset, Aleph-Alpha-GermanWeb, from three sources: (1) Common Crawl [13] web data not included in FineWeb2; (2) FineWeb2 (ODC-By v1.0); and (3) synthetic data conditioned on actual, organic web data. Our contributions are:
@@
    We trained two different model architectures, a 1 billion parameter Llama-style model [16] and a tokenizer-free 8 billion parameter model [35], on our dataset and FineWeb2. With these models, we evaluated a range of German language benchmarks, including Multilingual Massive Multitask Language Understanding (MMMLU) [19]. Our evaluations show our dataset outperforms FineWeb2 [21], even when combined with human-curated high-quality data sources such as Wikipedia at the 8 billion parameter scale. Our findings support the growing body of evidence that model-based data curation and synthetic data generation can significantly enhance pre-training datasets; and
@@
    We make Aleph-Alpha-GermanWeb publicly available to the research community at https://huggingface.co/datasets/Aleph-Alpha/Aleph-Alpha-GermanWeb, in the hope that it will contribute to advancements in German-language LLM pre-training.
@@ 2.1 Curating Common Crawl
We followed the RefinedWeb pipeline [39], as adapted for German in FineWeb2 [21], with some adjustments. We implemented all steps of the pipeline using NeMo Curator [23] (Apache-2.0). We summarise the number of web documents our pipeline yielded in Table 1.
@@
We downloaded six Common Crawl [13] (CC BY-SA 4.0) web data dumps, spanning from September 2024 to February 2025. Each dump captured raw web page data, e.g., HTML, metadata such as URLs, and text extracts.
@@
We used the URL filter developed in RefinedWeb [39] to filter out: (i) fraudulent and adult websites from a blocklist of 4.6M domains; (ii) additional URLs based on the presence of words associated with fraudulent or adult content; and (iii) content from so-called ‘high-quality’ websites such as Wikipedia, arXiv, etc., which allows these datasets to be mixed in at a later point and in controlled proportions.
@@
Natural language texts are embedded into raw web data via HTML and other web programming languages. Since we wish to focus on natural language, we extract the natural language from the raw web data using an extraction tool. RefinedWeb used the trafilatura text extractor [5]. However, subsequent work [25] constructing an English-language dataset from Common Crawl data showed that using the resiliparse text extractor [6] performed better on downstream LLM evaluations. We therefore use resiliparse for text extraction.
@@
RefinedWeb [39] used the fastText language classifier of CCNet [51] at the document level. This classifier was trained on character n-grams from Wikipedia and supports 176 natural languages. Using this classifier, we removed documents for which the top language score was not German.
@@
Web-scraped documents can include repetitive phrases or words. Heuristic methods for detection and removal of these repetitions were notably developed by Rae et al. [41]. We used these methods with the repetition identification and removal thresholds set in FineWeb2 [21] for German.
Repetition Type Threshold Measurement Unit
Duplicate whole lines $28.2\%$ of lines
Duplicate paragraphs $30.0\%$ of lines
Duplicate paragraphs $20.0\%$ of lines or paragraphs
Lines/paragraphs with many repeated characters $20.0\%$ of lines
Top 2-grams $7.7\%$ of characters
Top 3-grams $10.1\%$ of characters
Top 4-grams $12.3\%$ of characters
All duplicate 5-grams $14.2\%$ of characters
All duplicate 6-grams $12.7\%$ of characters
All duplicate 7-grams $11.5\%$ of characters
All duplicate 8-grams $10.6\%$ of characters
All duplicate 9-grams $9.7\%$ of characters
All duplicate 10-grams $8.8\%$ of characters
Table 2: Repetition removal thresholds for methods introduced in Rae et al. [41], with thresholds set as in FineWeb2 [21] for German.
@@
We utilized additional document-level heuristics introduced in Rae et al. [41] and adapted in FineWeb2 [21]; namely, we required documents to have:
  1. (i)
    >50 and <100,000 words;
  2. (ii)
    a mean word length of <14 characters;
  3. (iii)
    a symbols-to-words ratio <0.1, where symbols are “#” and “…”;
  4. (iv)
    <90% of lines starting with a bullet point;
  5. (v)
    <30% of lines ending with ellipses;
  6. (vi)
    >77.4% of words containing at least one alphabetic character; and
  7. (vii)
    at least two of the following list of common German words: “der”, “und”, “die”, “in”, “von”, “im”, “den”, “des”, “mit”, “das”, “er”, “dem”, “als”, “wurde”, “für”.
@@
Rae et al. [41] also introduced line-wise corrections to replace or remove unwanted text within documents. We opted instead for a stricter setting by removing documents if:
  1. (i)
    >15% of the document consists of numbers;
  2. (ii)
    >50% of the lines have >50% uppercase characters;
  3. (iii)
    the average words per line is <10 words; or
  4. (iv)
    >40% of paragraphs contain boilerplate strings, e.g., “terms of use”, “privacy policy”, etc..
@@
By first hashing the text of each document, we performed exact deduplication to remove all but one copy of identical documents, i.e., where there are two or more documents whose strings are equal in the dataset, we kept only one of these documents. We performed this deduplication procedure over the entire dataset. This approach is notably unlike FineWeb2 [21], which retained a certain number of duplicate documents according to their duplication rate within the overall dataset, a technique they refer to as ‘rehydration’.
@@
Our document-level fuzzy deduplication methodology closely follows Smith et al. [44], employing a multi-stage process to identify and remove near-duplicates within the corpus. The procedure began with computing MinHash signatures [8] using character-based 5-grams with 23-character sequences, where the number 23 was calculated by assuming an average of 4.5 characters per word. We then used locality sensitive hashing (LSH) [15] to identify candidate duplicates and sort them into 14 buckets, which each had eight unique hashing functions. LSH buckets were then converted to edges for the connected components algorithm. This algorithm identified document clusters with high similarity, producing groups of near-duplicate documents that can then be removed from the corpus. We performed this deduplication procedure over the entire dataset.
@@ 2.2 Synthetic data generation
Inspired by similar work for English-language web data [28, 46], we performed synthetic data generation conditioned on actual – what we also refer to as ‘organic’ – data from FineWeb2. We showcase the five prompt templates we used for synthetic data generation in Appendix B, which are designed to generate educational or basic rephrasings, summarizations, question-answer pairs, and lists of factual information. For all synthetic generation, we used the 12 billion parameter Mistral-Nemo-Instruct-2407111https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 (apache 2.0) LLM.
In all cases, if the entire document was longer than a specified character threshold, we segmented it into chunks of characters with lengths equal to or less than the threshold. This was done to prevent the organic document from potentially taking up all remaining context space in the prompt and to improve the quality of the LLM’s responses – prior studies [28, 46] also noted for English-language that if too much organic text is inserted it degrades the synthetic data quality. To avoid splitting the documents mid-word, -sentence, or -paragraph, and thereby potentially giving impractical or nonsensical inputs to the LLM, we segmented the document texts using text-splitter [^2] https://github.com/benbrandt/text-splitter. For a given prompt and document, we synthesized data for all semantic chunks until the organic document’s data was depleted. For each synthesised output, we also performed some post-processing using regular expressions and filtering to remove what we call ‘LLM artifacts’, such as prefixes like “Here’s the rephrased version:” and repetitions of the prompt or input text.
For a large number of FineWeb2 documents, we synthesized data based on one or more of our prompts. This resulted in many documents, especially long ones (which were segmented into more chunks), being represented in multiple ways in the final synthetic dataset. As mentioned in Su, Kong, Lin et al. [46], there is a non-negative (albeit diminishing) performance gain seen with exactly the same data during multi-epoch pre-training [33], therefore it is reasonable to expect that additional ‘synthetic epochs’ conditioned on the same underlying organic data is not harmful up to a similar limit. In the work of Muennighoff et al. [33], this limit was shown to be approximately four epochs. This is why we chose to limit our prompt templates to five, such that there would be no more than five possible synthetic outputs conditioned on the same organic input when the whole dataset was used for training.

### v3

@@ Abstract
Scaling data quantity is essential for large language models (LLMs), yet recent findings show that data quality can significantly boost performance and training efficiency. We introduce a German-language dataset curation pipeline that combines heuristic and model-based filtering techniques with synthetic data generation. We use our pipeline to create Aleph-Alpha-GermanWeb, a 628B-word German pre-training dataset composed of three subsets drawing from: (1) Common Crawl web data (organic subset; 78B words), (2) FineWeb2 (organic subset; 235B), and (3) synthetically-generated data conditioned on actual, organic web data (synthetic subset; 329B). We evaluate our dataset by pre-training both a 1B Llama-style model and an 8B tokeniser-free hierarchical autoregressive transformer (HAT) from scratch. A comparison on German-language benchmarks, including MMMLU, shows significant performance gains of Aleph-Alpha-GermanWeb over FineWeb2 alone. This advantage holds at the 8B scale even when FineWeb2 is enriched by human-curated high-quality data sources such as Wikipedia. Our findings support the growing body of evidence that model-based data curation and synthetic data generation can significantly enhance LLM pre-training datasets.
Aleph-Alpha-GermanWeb: Improving German-Language LLM Pre-Training with Model-Based Data Curation and
Synthetic Data Generation
Thomas F Burns          Letitia Parcalabescu          Stephan Wäldchen          Michael Barlow
Gregor Ziegltrum          Volker Stampa          Bastian Harren          Björn Deiseroththanks: Corresponding author bjoern.deiseroth@aleph-alpha-research.com
Aleph Alpha Research
@@ 1 Introduction
In recent years, ever-larger and more capable large language models (LLMs) have been released. This trend has led to the identification of power-law correlations between the loss value and the number of LLM parameters or pre-training dataset size Hoffmann et al. (2022); Su et al. (2024b). As a result, researchers have argued Hoffmann et al. (2022) that simply increasing the number of LLM parameters without also increasing the pre-training dataset size yields diminishing returns. Additionally, considering the cost of LLM inference Samsi et al. (2023) in production use-cases – which can far exceed the model’s initial training cost – practitioners may weigh the cost-benefit trade-off of replacing a larger model with a smaller one to minimise inference costs Irugalbandara et al. (2024). To compensate for the reduced model size, a sufficiently-large training data budget would be required De Vries (2023); Hassid et al. (2024). However, this approach poses significant challenges for domains and languages where data is scarce.
Many recent examples in both pre-training Marion et al. (2023); Su et al. (2024a) and post-training Nguyen and He (2025); Ye et al. (2025) of LLMs indicate that improvements to data quality can make training much more efficient. This is because higher-quality data can reduce the data quantity needed to reach or exceed the same model performance. In fact, conventional data scaling laws, which suggest that a small percentage increase in performance requires increasing the amount of training data by an order of magnitude, appear to be broken when data quality is improved Sorscher et al. (2023). As a result, improving data quality allows us to train LLMs much more economically. Indeed, a recent 1.7 billion parameter language model pre-trained on 11 trillion tokens achieved new state-of-the-art results for its size by focussing heavily on data quality Allal et al. (2025).
@@
  1. 1.
    removing syntactic (near) duplication, e.g., identical documents (or with slightly different wording), which are typically straightforward to identify and remove Lee et al. (2022);
  2. 2.
    removing semantic duplication, e.g., different summaries of the same book, which we expect to maintain some of for the purposes of syntactic variance, but should probably only keep in proportion to the level of importance or complexity of the underlying concept Abbas et al. (2023);
  3. 3.
    removing bad data, e.g., code that does not compile Di et al. (2024) or language which is ungrammatical Penedo et al. (2024), less our models learn to repeat these mistakes; and
  4. 4.
    augmenting or upsampling good data, e.g., adding high-quality synthetic data generated by LLMs Long et al. (2024) or intentionally duplicating high-quality subsets of the data Penedo et al. (2025).
Many curation methods originated as human-led and heuristic approaches, e.g., preselecting sources with a reputation for high quality data or manual annotation and filtering of datasets Penedo et al. (2023, 2024). More recently, however, model-based approaches, e.g., training annotation models to classify data into different quality buckets Li et al. (2024) or using LLM-based text quality filtering Su et al. (2024a), have been shown to outperform heuristic methods. Additionally augmenting datasets with synthetically-generated data conditioned on high-quality organic data sources not only improves quantity but has been shown to further improve quality Su et al. (2024a).
To date, model-based and synthetic approaches exist mostly for English-language text Li et al. (2024); Su et al. (2024a). One of the most commonly-used and largest heuristically-filtered non-English web datasets is FineWeb2 Penedo et al. (2025), which adapts prior heuristic methods shown to be successful for English web data Penedo et al. (2023, 2024). Recently, model-based methods have been applied to a subset of languages from FineWeb2 to score them for quality, finding as little as 15% of the original dataset can be used to achieve comparable performance on 1 billion parameter LLMs Messmer et al. (2025).
Nevertheless, there remains an overall quantity issue; FineWeb2, which includes data from over 1,000 languages, is <20% the size of the English-only FineWeb dataset Penedo et al. (2024) (8TB vs. 47TB in total disk size). A single language like German constitutes <1.4% of FineWeb2, and model-based approaches only reduce this quantity further Messmer et al. (2025). Meanwhile, synthetic approaches to improve the quantity, diversity, and quality of non-English language pre-training data have focussed on machine translation Wang et al. (2025), which comes with a host of socio-technical issues Moorkens et al. (2024), and may lack the distinct syntactic or semantic advantages offered by synthetic approaches for LLM pre-training Maini et al. (2024). Here we aim to fill these gaps by constructing a dataset for state-of-the-art German pre-training.
@@
In this paper we present a data curation pipeline tailored to German-language LLM pre-training. We use it to derive our own dataset, Aleph-Alpha-GermanWeb, a 642B-word corpus composed of three complementary subsets drawn from three distinct sources: (1) Common Crawl Foundation web data not included in FineWeb2 (78B words); (2) FineWeb2 (ODC-By v1.0) (235B words); and (3) synthetic data conditioned on actual, organic web data (329B words). Our contributions are:
@@
    We trained two different model architectures, a 1 billion parameter Llama-style model Grattafiori et al. (2024) and a tokeniser-free 8 billion parameter model Neitemeier et al. (2025b), on each subset of GermanWeb in isolation, without mixing data across subsets. For comparison, we trained the same architectures on FineWeb2 under identical training conditions. With these models, we evaluated a range of German language benchmarks, including Multilingual Massive Multitask Language Understanding (MMMLU) Hendrycks et al. (2021). Our evaluations show that all three subsets of our dataset (synthetic, filtered Common Crawl, and our filtering of Fineweb2) outperform FineWeb2 Penedo et al. (2025), even when the latter is combined with human-curated high-quality data sources such as Wikipedia at the 8 billion parameter scale. Our findings support the growing body of evidence that model-based data curation and synthetic data generation can significantly enhance pre-training datasets; and
@@
    We make GermanWeb publicly available to the research community [^1] https://huggingface.co/datasets/Aleph-Alpha/Aleph-Alpha-GermanWeb, in the hope that it will contribute to advancements in German-language LLM pre-training.
@@ 2.1 Curating Common Crawl
We followed the RefinedWeb pipeline Penedo et al. (2023), as adapted for German in FineWeb2 Penedo et al. (2025), with some adjustments. We implemented all steps of the pipeline using NeMo Curator Jennings et al. (Apache-2.0). We summarise the number of web documents our pipeline yielded in Table 1.
@@
We downloaded six Common Crawl Foundation (CC BY-SA 4.0) web data dumps, spanning from September 2024 to February 2025. Each dump captured raw web page data, e.g., HTML, metadata such as URLs, and text extracts.
@@
We used the URL filter developed in RefinedWeb Penedo et al. (2023) to filter out: (i) fraudulent and adult websites from a blocklist of 4.6M domains; (ii) additional URLs based on the presence of words associated with fraudulent or adult content; and (iii) content from so-called ‘high-quality’ websites such as Wikipedia, arXiv, etc., which allows these datasets to be mixed in at a later point and in controlled proportions.
@@
Natural language texts are embedded into raw web data via HTML and other web programming languages. Since we wish to focus on natural language, we extract the natural language from the raw web data using an extraction tool. RefinedWeb used the trafilatura text extractor Barbaresi (2021). However, subsequent work Li et al. (2024) constructing an English-language dataset from Common Crawl data showed that using the resiliparse text extractor Bevendorff et al. (2018) performed better on downstream LLM evaluations. We therefore use resiliparse for text extraction.
@@
RefinedWeb Penedo et al. (2023) used the fastText language classifier of CCNet Wenzek et al. (2020) at the document level. This classifier was trained on character n-grams from Wikipedia and supports 176 natural languages. Using this classifier, we removed documents for which the top language score was not German.
@@
Web-scraped documents can include repetitive phrases or words. Heuristic methods for detection and removal of these repetitions were notably developed by Rae et al. (2022). We used these methods with the repetition identification and removal thresholds set in FineWeb2 Penedo et al. (2025) for German.
Repetition Type Threshold Metric
Duplicate lines $28.2\%$ # of lines
Duplicate paras $30.0\%$ # of lines
Duplicate paras $20.0\%$ # of paras
Repeated chars $20.0\%$ # of lines
Top 2-grams $7.7\%$ # of chars
Top 3-grams $10.1\%$ # of chars
Top 4-grams $12.3\%$ # of chars
Duplicate 5-grams $14.2\%$ # of chars
Duplicate 6-grams $12.7\%$ # of chars
Duplicate 7-grams $11.5\%$ # of chars
Duplicate 8-grams $10.6\%$ # of chars
Duplicate 9-grams $9.7\%$ # of chars
Duplicate 10-grams $8.8\%$ # of chars
Table 2: Repetition removal thresholds for methods introduced in Rae et al. (2022), with thresholds set as in FineWeb2 Penedo et al. (2025) for German. In the table, ‘paras’ is short for ‘paragraphs’ and ‘chars’ is short ‘characters’.
@@
We used additional document-level heuristics introduced in Rae et al. (2022) and adapted in FineWeb2 Penedo et al. (2025); namely, we required documents to have: (i) >50 and <100,000 words; (ii) a mean word length of <14 characters; (iii) a symbols-to-words ratio <0.1, where symbols are “#” and “…”; (iv) <90% of lines starting with a bullet point; (v) <30% of lines ending with ellipses; (vi) >77.4% of words containing at least one alphabetic character; and (vii) at least two of the following list of common German words: “der”, “und”, “die”, “in”, “von”, “im”, “den”, “des”, “mit”, “das”, “er”, “dem”, “als”, “wurde”, “für” (English translations: “the”, “and”, “the”, “in”, “of”, “in”, “the”, “of”, “with”, “that”, “he”, “the”, “as”, “was”, “for”).
@@
Rae et al. (2022) also introduced line-wise corrections to replace or remove unwanted text within documents. We opted instead for a stricter setting by removing documents if: (i) >15% of the document consists of numbers; (ii) >50% of the lines have >50% uppercase characters; (iii) the average words per line is <10 words; or (iv) >40% of paragraphs contain boilerplate strings, e.g., “terms of use”, “privacy policy”, etc..
@@
By first hashing the text of each document, we performed exact deduplication to remove all but one copy of identical documents, i.e., where there are two or more documents whose strings are equal in the dataset, we kept only one of these documents. We performed this deduplication procedure over the entire dataset. This approach is notably unlike FineWeb2 Penedo et al. (2025), which retained a certain number of duplicate documents according to their duplication rate within the overall dataset, a technique they refer to as ‘rehydration’.
@@
Our document-level fuzzy deduplication methodology closely follows Smith et al. (2022), employing a multi-stage process to identify and remove near-duplicates within the corpus. The procedure began with computing MinHash signatures Broder (1997) using character-based 5-grams with 23-character sequences, where the number 23 was calculated by assuming an average of 4.5 characters per word. We then used locality sensitive hashing (LSH) Gionis et al. (1999) to identify candidate duplicates and sort them into 14 buckets, which each had eight unique hashing functions. LSH buckets were then converted to edges for the connected components algorithm. This algorithm identified document clusters with high similarity, producing groups of near-duplicate documents that can then be removed from the corpus. We performed this deduplication procedure over the entire dataset.
@@ 2.2 Synthetic data generation
Inspired by similar work for English-language web data Maini et al. (2024); Su et al. (2024a), we performed synthetic data generation conditioned on actual – what we also refer to as ‘organic’ – data from FineWeb2. We showcase the five prompt templates we used for synthetic data generation in Appendix B, designed to generate educational or basic rephrasings, summarisations, question-answer pairs, and lists of factual information. For all synthetic generation, we used the 12 billion parameter Mistral-Nemo-Instruct-2407222https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 (Apache 2.0) LLM.
In all cases, if the entire document was longer than a specified character threshold, we segmented it into chunks of characters with lengths equal to or less than the threshold. This was done to prevent the organic document from potentially taking up all remaining context space in the prompt and to improve the quality of the LLM’s responses – prior studies Maini et al. (2024); Su et al. (2024a) also noted for English-language that if too much organic text is inserted it degrades the synthetic data quality. To avoid splitting the documents mid-word, -sentence, or -paragraph, and thereby potentially giving impractical or nonsensical inputs to the LLM, we segmented the document texts using text-splitter [^3] https://github.com/benbrandt/text-splitter. For a given prompt and document, we synthesized data for all semantic chunks until the organic document’s data was depleted. For each synthesised output, we also performed some post-processing using regular expressions and filtering to remove what we call ‘LLM artifacts’, such as prefixes like “Here’s the rephrased version:” and repetitions of the prompt or input text.
For a large number of FineWeb2 documents, we synthesized data based on one or more of our prompts. This resulted in many documents, especially long ones (which were segmented into more chunks), being represented in multiple ways in the final synthetic dataset. As mentioned in Su et al. (2024a), there is a non-negative (albeit diminishing) performance gain seen with exactly the same data during multi-epoch pre-training Muennighoff et al. (2023), therefore it is reasonable to expect that additional ‘synthetic epochs’ conditioned on the same underlying organic data is not harmful up to a similar limit. In the work of Muennighoff et al. (2023), this limit was shown to be approximately four epochs. We thus chose to limit our prompt templates to five, such that there would be no more than five possible synthetic outputs conditioned on the same organic input when the whole dataset was used for training.